# Geometry-Informed Neural Networks

### v3

@@ 4.3 Engineering design case study
We further solve variations of the above problem to highlight the user’s ability to tune the problem and the resulting solutions. In particular, we include an additional topology constraint on the number of holes and employ another surface smoothness $E_{\text{log}}$. The variety of produced shapes is illustrated in Figure 1. This also illustrates the robustness of ALM and GINNs to problem variations.
Multiple solutions.
Upon the addition of the diversity constraint, GINNs do not only produce multiple solutions, but we also observe the emergence of a latent space structure. This is illustrated in Figure 6 using a $\dim(z)=2$ latent space from which $k=9$ random samples are drawn every training iteration. Traversing the latent space of the trained GINN produces continuously morphing feasible shapes, i.e. the model generalizes. Furthermore, the latent space is organized. However, we find that the learned structure depends on the exact setup of the diversity constraint. In particular, we observe a more pronounced organization emerge for larger $\delta_{\text{min}}$ while over-specifying it impacts convergence.
@@
The convergence behavior of the aforementioned runs with and without diversity is shown in Figures 9 and 10. Despite having up to seven loss terms, adaptive ALM automatically balances these and minimizes each constraint violation. However, the variance in several losses remains high. This is largely due to the diversity and smoothness terms, which are hard to optimize and increase the necessary number of iterations by roughly a factor of two and five, respectively. In addition, the necessary surface point sampling accounts for roughly 50% of the runtime, motivating improved strategies. Overall, training a single shape takes roughly 10K iterations and 1 hour. Similarly, the discussed diverse model takes 50K iterations and 72 hours.
Surface smoothness.
As illustrated in Figure 3, the smoothness term significantly improves the visual quality of the shapes. However, it is a difficult objective. First, the values of the surface-strain $E$ span several orders of magnitude and small deviations in the surface sampling cause large variance. Hence, the strategies described in Section 4.1 are critical to stabilize training. Employing $E_{\text{log}}$ improves convergence further, but also leads to a different surface quality due to the concavity property of $\log{}$ (cf. top and bottom rows in Figure 1). In previous works optimizing integrals over fixed domains, adaptive ALM deals with the range issue by adjusting location-specific penalties. However, this strategy is not directly applicable to our moving surfaces, which presents a unique future challenge. In addition, curvature is a second-order differential operator and is expected to be ill-conditioned, motivating the use of second-order optimizers to further refine the results [14, 13].
Figure 6: With diversity, GINNs do not only produce multiple solutions but also learn a latent space structure. Traversing the 2D latent space continuously morphs solutions, i.e., the model generalizes. The latent space is also organized – a central bulky shape becomes thinner in the radial direction and the axes can be identified by how the shape connects on the sides. Figure 8 shows a $9\times 9$ version.
@@ 5 Conclusion
We have introduced geometry-informed neural networks demonstrating shape-generative modeling driven solely by geometric constraints and objectives. After formulating the learning problem and discussing key theoretical and practical aspects, we applied GINNs to several validation problems and a realistic engineering task.
@@
GINNs combine several known and novel components, each of which warrants an in-depth study of theoretical and practical aspects, including alternative shape distances and their aggregation into a diversity, conditioning mechanisms, and a broader range of constraints.
In this work, we focused on building the conceptual framework of GINNs and validating it experimentally. This included a realistic generative design task. However, we considered a modified version of the original task and did not compare to established topology-optimization methods as this required an integration of a PDE solver – a task that future work should address.
Even though ALM is a significant improvement over the naive approach of manually weighted loss terms, the recent literature on multi-objective and second-order optimizers suggest further possible improvements.
@@ Acknowledgments
The ELLIS Unit Linz, the LIT AI Lab, and the Institute for Machine Learning are supported by the Federal State of Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank the European High Performance Computing initiative for providing computational resources (EHPC-DEV-2023D08-019, 2024D06-055, 2024D08-061). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF, and the NVIDIA Corporation.
@@
Appendix A Implementation and experimental details
@@
We report additional details on the experiments and their implementation. We run all experiments on a single GPU (one of NVIDIA RTX2080Ti, RTX3090, A40, or P40). For single-shape training, the maximum GPU memory requirements are ca. 9GB for the jet engine bracket and less than a GB for the rest. For multi-shape training the maximum GPU memory requirements are ca. 45GB for the jet engine bracket (9 shapes) and ca. 7GB for the obstacle problem (16 shapes).
A.1 Reaction-diffusion
See Appendix B.
A.2 Plateau’s problem
@@
The model is an MLP with $[3,256,256,256,1]$ neurons per layer and the $\operatorname{tanh}$ activation. We train with Adam (default parameters) for 10000 epochs with a learning rate of $10^{-3}$ taking around three minutes. The three losses (interface, mean curvature, and eikonal) are weighted equally but mean curvature loss is introduced only after 1000 epochs. To facilitate a higher level of detail, the corner points of the prescribed interface are weighted higher.
A.3 Parabolic mirror
@@
The model is an MLP with $[2,40,40,1]$ neurons per layer and the $\operatorname{tanh}$ activation. We train with Adam (default parameters) for 3000 epochs with a learning rate of $10^{-3}$ taking around ten seconds.
A.4 Obstacle
@@
Consider the domain $\mathcal{X}=[-1,1]\times[-0.5,0.5]$ and the design region that is a smaller rectangular domain with a circular obstacle in the middle: $\mathcal{E}=([-0.9,0.9]\times[-0.4,0.4])\setminus\{x_{1}^{2}+x_{2}^{2}\leq 0.1^{2}\}$. There is an interface consisting of two vertical line segments ={(±$\mathcal{I}=\{(\pm 0.9,x_{2})|-0.4\leqslant x_{2}\leqslant 0.4\}$ with the prescribed outward facing normals $\bar{n}(\pm 0.9,-0.4\leqslant x_{2}\leqslant 0.4)=(\pm 1,0)$.
@@
The neural network model $f$ should be at least twice differentiable with respect to the inputs $x$, as necessitated by the computation of surface normals and curvatures. Since the second derivatives of an ReLU MLP is zero everywhere, we use the softplus activation function as a simple baseline. In addition, we add residual connections [66] to mitigate the vanishing gradient problem and facilitate learning. We denote this architecture with "softplus-MLP". We train a softplus-MLP with Adam (default settings) and the hyperparameters in Table 2.
@@
For training the conditional models, we approximate the one-dimensional latent set $Z=[-1,1]$ with $N=16$ fixed equally spaced samples. This enables the reuse of some calculations across epochs and results in a well-structured latent space, illustrated through latent space interpolation in Figure 5.
Hyperparameter Obstacle (2D) JEB (3D)
Architecture Residual-MLP WIRE
Hidden layers $[512,512,512,512]$ $[128,128,128]$
Activation softplus Gabor wavelet
$\omega_{0}$ for WIRE n/a 18.0
$s_{0}$ for WIRE n/a 6.0
Learning rate 0.001 0.001
Learning rate schedule $0.5^{t/1000}$ $0.5^{t/5000}$
Iterations 3000 10000
Table 2: Hyperparameters for the generative 2D obstacle and 3D jet engine bracket experiments. The input is a 2D or 3D point concatenated with a 1D latent vector. For both experiments, the initial learning rate is halved every 1000 (Obstacle) or 5000 (jet engine bracket) iterations. The hidden layers do not include an input layer of input-dimension $\dim(x)+\dim(z)$, whereas $x$ is a coordinate and $z$ is a modulation vector, and an output layer of output-dimension 1. Interestingly, the WIRE network overall had fewer parameters, while fitting a more complex shape.
@@
The total training wall-clock time is around 10 minutes for a single shape and approximately 60 minutes for 16 shapes. These numbers are without applying a smoothness loss.
A.5 Jet engine bracket
The jet engine bracket (JEB) is our most complex experiment. We tested different architectures (c.f. Figure 7) and found that WIRE [34] produced the best results, while being easier to train with the augmented Lagrangian method than softplus-MLP or SIREN [20]. We train with Adam (default settings) and the hyperparameters summarized in Table 2. To decrease the training time, we use multi-processing to asynchronously create diagnostic plots or computing the PH loss for multiple shapes.
@@ A.5 Jet engine bracket
For the jet engine bracket settings, early experiments indicated that the softplus-MLP cannot satisfy the given constraints. We therefore employ a WIRE network [34], which is biased towards higher frequencies of the input signal. As mentioned by the authors, the spectral properties of a WIRE model are relatively robust. Several values for $\omega_{0}$ and $s_{0}$, which control the frequency and scale of the gaussian of the first layer at initialization, were tested. As there was no big difference in the results, we fixed them to $\omega_{0}=18$ and $s_{0}=6$ For more detailed results, we refer to Section C.
@@
Figure 8: GINNs produce diverse shapes with a structured latent space. The shapes morph continiously into one another when traversing the 2 dimensional latent space. These shapes are produced by the same model as Figure 6. A trained GINN allows the user to sample densely in the latent space with shapes all meeting the constraints: Interfaces are modeled correctly, shapes are not disconnected or leave the design space.
@@
The curse of dimensionality implies that with higher dimensions, exponentially (in the number of dimensions) more points are needed to cover the space equidistantly. Therefore, in 3D, substantially more points (and consequently memory and compute) are needed than in 2D. In our experiments, we observe that a low spatial resolution around the interfaces prevents the model from learning high-frequency details, likely due to a stochastic gradient. Increased spatial resolution results in a better learning signal and the model picks up the details easier. To facilitate learning we additionally increase the resolution around the interfaces.
Appendix B Generative PINNs
Having developed a generative GINN that is capable of producing diverse solutions to an under-determined problem, we ask if this idea generalizes to other areas. In physics, problems are often well-defined and have a unique solution. However, cases exist where the initial conditions are irrelevant and a non-particular PDE solution is sufficient, such as in chaotic systems or animations.
We conclude the experimental section by demonstrating an analogous concept of generative PINNs on a reaction-diffusion system. Such systems were introduced by Turing [67] to explain how patterns in nature, such as stripes and spots, can form as a result of a simple physical process of reaction and diffusion of two substances. A celebrated model of such a system is the Gray-Scott model [68], which produces a variety of patterns by changing just two parameters – the feed-rate $\alpha$ and the kill-rate $\beta$ – in the following PDE:
This PDE describes the concentration $u,v$ of two substances $U,V$ undergoing the chemical reaction $U+2V\rightarrow 3V$. The rate of this reaction is described by $uv^{2}$, while the rate of adding $U$ and removing $V$ is controlled by the parameters $\alpha$ and $\beta$. Crucially, both substances undergo diffusion (controlled by the coefficients $D_{u},D_{v}$) which produces an instability leading to rich patterns around the bifurcation line $\alpha=4(\alpha+\beta)^{2}$.
Computationally, these patterns are typically obtained by evolving a given initial condition $u(x,t=0)=u_{0}(x)$, $v(x,t=0)=v_{0}(x)$ on some domain with periodic boundary conditions. A variety of numerical solvers can be applied, but previous PINN attempts fail without data [69]. To demonstrate a generative PINN on a problem that admits multiple solutions, we omit the initial condition and instead consider stationary solutions, which are known to exist for some parameters $\alpha,\beta$ [70]. We use the corresponding stationary PDE ($\partial u/\partial t=\partial v/\partial t=0$) to formulate the residual losses:
To avoid trivial (i.e. uniform) solutions, we encourage non-zero gradient with a loss term $-\max(1,\int_{\mathcal{D}}(\nabla u(x))^{2}+(\nabla v(x))^{2}\operatorname{d}\!{x})$. We find that architecture and initialization are critical (see Appendix B.1). Using the diffusion coefficients $D_{v}=1.2\times 10^{-5}$, $D_{u}=2D_{v}$ and the feed and kill-rates $\alpha=0.028,\beta=0.057$, the generative PINN produces diverse and smoothly changing pattern of worms, illustrated in Figure 4. To the best of our knowledge, this is the first PINN that produces 2D Turing patterns in a data-free setting.
B.1 Experimental details
We use two identical SIREN networks for each of the fields $u$ and $v$. They have two hidden layers of widths 256 and 128. We enforce periodic boundary conditions on the unit domain $\mathcal{X}=[0,1]^{2}$ through the encoding $x_{i}\mapsto(\sin{2\pi x_{i}},\cos{2\pi x_{i}})$ for $i=1,2$. With this encoding, we use $\omega_{0}=3.0$ to initialize SIREN. We also find that the same shaped Fourier-feature network [33] with an appropriate initialization of $\sigma=3$ works equally well.
Appendix C Evaluation

### v4

@@ 4.3 Discussion
The variations of the bracket problem illustrated in Figure 1 highlight the user’s ability to tune the problem and the resulting solutions. We solve these with the same setup and hyperparameters, illustrating the robustness of GINNs and the adaptive ALM approach to constrained optimization.
@@
A representative convergence behavior of the training is shown in Figures 11 and 12. Despite up to seven loss terms, adaptive ALM automatically balances these and minimizes each constraint violation. However, the variance in several losses remains high. This is largely due to the diversity and smoothness terms, which are hard to optimize and increase the necessary number of iterations by roughly a factor of two and five, respectively. Curvature is a second-order differential operator and is expected to be ill-conditioned, motivating the future use of second-order optimizers (Ryck et al., 2024; Rathore et al., 2024).
Runtime
of a single bracket shape is roughly 10K iterations or 30 minutes. Similarly, the diverse model trains for around 50K iterations or 5 hours on a single GPU. Of the total time, the surface strain takes roughly 10% (due to the Hessian) and the PH solver 75% (expensive multi-processed CPU task). The runtime also increases when ablating the eikonal constraint as it destroys the geometric regularity of the implicit function, hindering efficient surface point sampling that usually takes 15% of total time. More details are available in Appendix C.3.
Ablations
are performed in Appendix C.2, measuring the expected role of each constraint, as well as ALM. Less obvious is the strong impact of the eikonal constraint on the connectedness, smoothness, and diversity metrics, again due to the lack of geometric regularity of the implicit function. A progressive ablation of all requirements of the bracket problem is illustrated in Figure 3. Figure 8 also visually compares different NF models, highlighting the superiority of WIRE over softplus-MLP and SIREN due to their spectral behavior.
Figure 6: With the diversity constraint, GINNs not only produce multiple solutions but also discover a latent space structure. Traversing the 2D latent space morphs solutions, i.e., the model generalizes. The latent space is also organized – a central bulky shape becomes thinner in the radial direction, and the axes can be identified by how the shape connects on the sides. Figure 9 shows a large $9\times 9$ version.
@@ 5 Conclusion
We have introduced geometry-informed neural networks demonstrating shape-generative modeling driven solely by geometric constraints and objectives. After formulating the learning problem and discussing key theoretical and practical aspects, we applied GINNs to a range of problems.
@@
GINNs combine several known and novel components, each of which warrants an in-depth study of theoretical and practical aspects, including alternative shape distances, their aggregation into diversity, conditioning mechanisms, constraints, and optimization.
In this work, we focused on building the conceptual framework of GINNs and validating it experimentally. This included a realistic engineering design task. However, we considered a modified version of the original task and did not compare to established topology-optimization methods as this required the integration of a PDE solver – a task we address in a follow-up work.
Even though ALM is a significant improvement over the naive approach of manually weighted loss terms, the recent literature on multi-objective and second-order optimizers suggests further possible improvements.
@@ Acknowledgments
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We thank the European High Performance Computing initiative for providing computational resources (EHPC-DEV-2023D08-019, 2024D06-055, 2024D08-061). We thank NXAI GmbH, Audi AG, Silicon Austria Labs (SAL), Merck Healthcare KGaA, GLS (Univ. Waterloo), TÜV Holding GmbH, Software Competence Center Hagenberg GmbH, dSPACE GmbH, TRUMPF SE + Co. KG.
@@
Impact Statement
Our work aims to advance the field of machine learning and may contribute to its broader societal impact. In addition, there is an ongoing discussion on the rights to data, since data is fundamental to training the majority of current machine learning models. The demonstrated data-free approach to generative modeling brings forth a less explored perspective on this matter. It circumvents the copyright problem and facilitates practitioners who lack exclusive access to datasets. However, for the foreseeable future, the applicability and hence the impact is limited to scientific and engineering applications. The demonstrated results, in particular, might foster a path toward improved approaches to engineering design.
Appendix A Problem-Constraint Matrix
Table 2 provides an overview of the variety of considered tasks using a problem-constraint matrix. Not captured by this matrix are also the differences in domain dimension, shape representation, and problem symmetries.
Table 2: Problem-constraint matrix.
Turing patterns Plateau’s problem Parabolic mirror Obstacle Wheels Jet engine bracket
Reaction-diffusion PDE +
Interface + + + + +
Mean-curvature +
Reflection +
Design region + + +
Prescribed normal + + +
Connectedness + + +
Diversity + + + +
Eikonal + + +
Rotational symmetry +
Holes +
Surface strain +
Minimum thickness + +
Appendix B Implementation and Experimental Details
@@
We report additional details on the experiments and their implementation. We run all experiments on a single GPU (one of NVIDIA RTX2080Ti, RTX3090, A40, P40, or A100-SXM). The largest GPU memory requirement is 16GB for the multi-shape training of the jet engine bracket (9 shapes).
Surface sampling
is required to estimate the surface integrals for smoothness and diversity. We first sample points in the envelope and project them onto the surface using Newton iterations. We then repel the points on the surface to achieve a more uniform distribution similar to Yifan et al. (2020). Finally, we exclude points sampled within a small distance to the interface $\mathcal{I}$, as the surface should not change here. We also begin to sample surface points and compute the surface integrals only after a warm-up phase of 500 iterations. In combination, these aspects lead to a lower variance and better convergence.
B.1 Reaction-Diffusion
The high-level idea of GINNs is to find diverse solutions to an under-determined problem. While the main focus of the paper is on geometry, we show that this idea generalizes to other areas. In physics, problems are often well-defined and have a unique solution. However, cases exist where the initial conditions are irrelevant and a non-particular PDE solution is sufficient, such as in chaotic systems or animations.
We demonstrate an analogous concept of generative PINNs on a reaction-diffusion system. Such systems were introduced by Turing (1952) to explain how patterns in nature, such as stripes and spots, can form as a result of a simple physical process of reaction and diffusion of two substances. A celebrated model of such a system is the Gray-Scott model (Pearson, 1993), which produces a variety of patterns by changing just two parameters – the feed-rate $\alpha$ and the kill-rate $\beta$ – in the following PDE:
This PDE describes the concentration $u,v$ of two substances $U,V$ undergoing the chemical reaction $U+2V\rightarrow 3V$. The rate of this reaction is described by $uv^{2}$, while the rate of adding $U$ and removing $V$ is controlled by the parameters $\alpha$ and $\beta$. Crucially, both substances undergo diffusion (controlled by the coefficients $D_{u},D_{v}$), which produces an instability leading to rich patterns around the bifurcation line $\alpha=4(\alpha+\beta)^{2}$.
Computationally, these patterns are typically obtained by evolving a given initial condition $u(x,t=0)=u_{0}(x)$, $v(x,t=0)=v_{0}(x)$ on some domain with periodic boundary conditions. A variety of numerical solvers can be applied, but previous PINN attempts fail without data (Giampaolo et al., 2022). To demonstrate a generative PINN on a problem that admits multiple solutions, we omit the initial condition and instead consider stationary solutions, which are known to exist for some parameters $\alpha,\beta$ (McGough & Riley, 2004). We use the corresponding stationary PDE ($\partial u/\partial t=\partial v/\partial t=0$) to formulate the residual losses:
To avoid trivial (i.e. uniform) solutions, we encourage non-zero gradient with a loss term $-\max(1,\int_{\mathcal{D}}(\nabla u(x))^{2}+(\nabla v(x))^{2}\operatorname{d}\!{x})$. We find that architecture and initialization are critical (described below). Using the diffusion coefficients $D_{v}=1.2\times 10^{-5}$, $D_{u}=2D_{v}$ and the feed and kill-rates $\alpha=0.028,\beta=0.057$, the generative PINN produces diverse and smoothly changing pattern of worms, illustrated in Figure 4. To the best of our knowledge, this is the first PINN that produces 2D Turing patterns in a data-free setting.
Experimental details.
We use two identical SIREN networks for each of the fields $u$ and $v$. They have two hidden layers of widths 256 and 128. We enforce periodic boundary conditions on the unit domain $\mathcal{X}=[0,1]^{2}$ through the encoding $x_{i}\mapsto(\sin{2\pi x_{i}},\cos{2\pi x_{i}})$ for $i=1,2$. With this encoding, we use $\omega_{0}=3.0$ to initialize SIREN. We also find that the same shaped Fourier-feature network (Tancik et al., 2020) with an appropriate initialization of $\sigma=3$ works equally well.
B.2 Plateau’s Problem
@@
The model is an MLP with $[3,256,256,256,1]$ neurons per layer and the $\operatorname{tanh}$ activation. We train with Adam (default parameters) for 10,000 epochs with a learning rate of $10^{-3}$, taking around three minutes. The three losses (interface, mean curvature, and eikonal) are weighted equally, but the mean curvature loss is introduced only after 1000 epochs. To facilitate a higher level of detail, the corner points of the prescribed interface are weighted higher.
B.3 Parabolic Mirror
@@
The model is an MLP with $[2,40,40,1]$ neurons per layer and the $\operatorname{tanh}$ activation. We train with Adam (default parameters) for 3000 epochs with a learning rate of $10^{-3}$, taking around ten seconds.
B.4 Obstacle
@@
Consider the domain $\mathcal{X}=[-1,1]\times[-0.5,0.5]$ and the design region that is a smaller rectangular domain with a circular obstacle in the middle: $\mathcal{E}=([-0.9,0.9]\times[-0.4,0.4])\setminus\{x_{1}^{2}+x_{2}^{2}\leq 0.1^{2}\}$. There is an interface consisting of two vertical line segments ={(±$\mathcal{I}=\{(\pm 0.9,x_{2})|-0.4\leq x_{2}\leq 0.4\}$ with the prescribed outward facing normals $\bar{n}(\pm 0.9,-0.4\leq x_{2}\leq 0.4)=(\pm 1,0)$.
@@
The neural network model $f$ should be at least twice differentiable with respect to the inputs $x$, as necessitated by the computation of surface normals and curvatures. Since the second derivatives of an ReLU MLP are zero everywhere, we use the softplus activation function as a simple baseline. In addition, we add residual connections (Dugas et al., 2000) to mitigate the vanishing gradient problem and facilitate learning. We denote this architecture with ”softplus-MLP”. We train a softplus-MLP with $4\times 512$ hidden layers with Adam (default settings).
@@
For training the conditional models, we approximate the one-dimensional latent set $Z=[-1,1]$ with $N=16$ fixed equally spaced samples. This enables the reuse of some calculations across epochs and results in a well-structured latent space, illustrated through latent space interpolation in Figure 5(c).
@@
The total training wall-clock time is around 10 minutes for a single shape and approximately 60 minutes for 16 shapes.
B.5 Wheel
The problem setup is as described in the main text. The ML setup is shared with the bracket problem. A larger illustration of the results is available in Figure 7, showing diverse shapes and a structured latent space.
Figure 7: GINNs produce diverse shapes with a structured latent space. The shapes morph continuously into one another when traversing the 2D latent space. From the top-left to bottom-right, the holes become smaller, while from bottom-left to top-right, the holes move inwards.
B.6 Jet Engine Bracket
The jet engine bracket (JEB) is our most complex experiment. We tested different architectures (c.f. Figure 8) and found that WIRE (Saragadam et al., 2023) produced the best results, while being easier to train with ALM than softplus-MLP or SIREN (Sitzmann et al., 2020). We train the WIRE with $3\times 128$ hidden layers with Adam (default settings) and a learning rate scheduler $0.5^{t/10000}$ for $t=10000$ iterations for the single shape and $t=50000$ for multiple shapes. To decrease the training time, we use multi-processing to asynchronously create diagnostic plots or computing the PH loss for multiple shapes.
@@ B.6 Jet Engine Bracket
For the jet engine bracket settings, early experiments indicated that the softplus-MLP cannot satisfy the given constraints. We therefore employ a WIRE network (Saragadam et al., 2023), which is biased towards higher frequencies of the input signal. As mentioned by the authors, the spectral properties of a WIRE model are relatively robust. Several values for $\omega_{0}$ and $s_{0}$, which control the frequency and scale of the Gaussian of the first layer at initialization, were tested. As there was no big difference in the results, we fixed them to $\omega_{0}=18$ and $s_{0}=6$ For more detailed results, we refer to Section C.
@@
Figure 9: GINNs produce diverse shapes with a structured latent space. The shapes morph continuously into one another when traversing the 2-dimensional latent space. These shapes are produced by the same model as Figure 6. A trained GINN allows the user to sample densely in the latent space with shapes all meeting the constraints: Interfaces are modeled correctly, shapes are not disconnected or leave the design space.
@@
The curse of dimensionality implies that with higher dimensions, exponentially (in the number of dimensions) more points are needed to cover the space equidistantly. Therefore, in 3D, substantially more points (and consequently memory and compute) are needed than in 2D. In our experiments, we observe that a low spatial resolution around the interfaces prevents the model from learning high-frequency details, likely due to a stochastic gradient. Increased spatial resolution results in a better learning signal, and the model picks up the details more easily. To facilitate learning, we additionally increase the resolution around the interfaces.
Appendix C Additional Evaluation
After defining the metrics used in our experiments in Appendix C.1, we perform an additional study in Appendix C.2 ablating each constraint, as well as the ALM, for three tasks. Additionally, in Appendix C.3, we measure the impact of each constraint on the runtime. Finally, Appendix C.4 discusses and applies two baseline methods to the jet engine bracket problem.